# A Weibull Mixture Cure Frailty Model for High-dimensional Covariates

### v1

@@ 3 Inference in the Penalized Mixture Cure Frailty Model
$L(\alpha,\gamma,\theta,\bm{\beta},\mathbf{b}|\mathbf{D})$ $=$ $\prod_{i=1}^{n}\left\{f_{pop}(t_{i}|\mathbf{x}_{i}\mathbf{,z}_{i})\right\}^{\delta_{i}}\left\{S_{pop}(t_{i}|\mathbf{x}_{i}\mathbf{,z}_{i})\right\}^{1-\delta_{i}}$
$=$ $\prod_{i=1}^{n}\left\{\frac{e^{\mathbf{bz}_{i}}}{1+e^{\mathbf{bz}_{i}}}\alpha\gamma t_{i}^{\gamma-1}e^{\bm{\beta}\mathbf{x}_{i}}\left(1+\frac{\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}}{\theta}\right)^{-\theta-1}\right\}^{\delta_{i}}$
$\left\{\frac{1}{1+e^{\mathbf{bz}_{i}}}\left(1+e^{\mathbf{bz}_{i}}\left(1+\frac{\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}}{\theta}\right)^{-\theta}\right)\right\}^{1-\delta_{i}},$
where $\alpha,\gamma$ are the unknown parameters of the Weibull distribution, $\theta$ is the unknown parameter of the gamma distribution used for the frailty, $\mathbf{b}\in\mathbb{R}^{P_{1}+1}$ and $\bm{\beta}\in\mathbb{R}^{P_{2}}$ are vectors of unknown regression coefficients for the covariates $\mathbf{x}$ and $\mathbf{z}$, respectively, where we indicate $\mathbf{bz}_{i}=b_{0}+\mathbf{b}_{u}\mathbf{z}_{u,i}+\mathbf{b}_{p}\mathbf{z}_{p,i}$ and $\bm{\beta}\mathbf{x}_{i}=\bm{\beta}_{u}\mathbf{x}_{u,i}+\bm{\beta}_{p}\mathbf{x}_{p,i}$, $i=1,...,n$, . The maximum likelihood estimator of the unknown parameters can be obtained by direct maximization of the observed likelihood function in (3) when the number of unknown parameters is not large. Since we would like to consider high-dimensional covariates and perform variable selection via penalization, direct maximization of the likelihood function will not provide parameter estimates. We therefore propose to adapt the E-M algorithm to include regularization methods to the purposes of parameter estimation and variable selection in penMCFM. Details on the proposed implementation of the EM for penMCFM are given in Section 3.2, after recalling the standard EM for MCFM in Section 3.1. For the sake of comparisons in the simulation studies, we have also adapted the GMIFS method from [38] to obtain parameter estimations in penMCFM using the observed likelihood function in (3), and details on the implementation of this method are given in Section 3.4.
@@ 3.1 EM Algorithm for the standard MCFM
$S_{pop}(t_{i}|\mathbf{x}_{i}\mathbf{,z}_{i}\mathbf{,}w_{i})=1-\pi(\mathbf{z}_{i})+\pi(\mathbf{z}_{i})\exp\left(-w_{i}\alpha t_{i}^{\gamma}e^{\bm{\beta x}_{i}}\right),$
@@
$S(t_{i}|\mathbf{x}_{i}\mathbf{,}w_{i},Y_{i}\mathbf{=}1)=\exp\left(-w_{i}\alpha t_{i}^{\gamma}e^{\bm{\beta x}_{i}}\right),\quad\text{and}\quad h(t_{i}|\mathbf{x,}w_{i},Y_{i}\mathbf{=}1)=w_{i}\alpha\gamma t_{i}^{\gamma-1}e^{\bm{\beta x}_{i}}.$
@@
$\pi(\mathbf{z}_{i})S(t_{i}|\mathbf{x}_{i}\mathbf{,}w_{i},Y_{i}\mathbf{=}1)\left[h(t_{i}|\mathbf{x,}w_{i},Y_{i}\mathbf{=}1)\right]^{\delta_{i}}$
@@
$l_{c_{2}}(\alpha,\gamma,\bm{\beta})=\sum_{i=1}^{n}\left[\delta_{i}\log\left(\alpha\gamma t_{i}^{\gamma-1}e^{\bm{\beta}\mathbf{x}_{i}}\right)-w_{i}y_{i}\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}\right],$
@@
$p_{i}^{(m)}\equiv E\left(y_{i}|\bm{\Lambda}^{(m)},\mathbf{D}\right)=\left.\delta_{i}+(1-\delta_{i})\frac{\pi(\mathbf{z}_{i})\left(1+\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}/\theta\right)^{-\theta}}{(1-\pi(\mathbf{z}_{i}))+\pi(\mathbf{z}_{i})\left(1+\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}/\theta\right)^{-\theta}}\right|_{\left(\bm{\Lambda}^{(m)},\mathbf{D}\right)},$ (8)
$a_{i}^{(m)}\equiv E\left(w_{i}|\bm{\Lambda}^{(m)},\mathbf{D}\right)=\left.\frac{\delta_{i}+\theta}{\theta+\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}}p_{i}+\frac{\delta_{i}+\theta}{\theta}(1-p_{i})\right|_{\left(\bm{\Lambda}^{(m)},\mathbf{D}\right)},$ (9)
$b_{i}^{(m)}\equiv E\left(\log w_{i}|\bm{\Lambda}^{(m)},\mathbf{D}\right)=\left.\left[\varphi(\delta_{i}+\theta)-\log(\theta+\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}})\right]p_{i}+\left[\varphi(\delta_{i}+\theta)-\log\theta\right](1-p_{i})\right|_{\left(\bm{\Lambda}^{(m)},\mathbf{D}\right)},$ (10)
$c_{i}^{(m)}\equiv E\left(w_{i}y_{i}|\Lambda^{(m)},\mathbf{D}\right)=\left.\frac{\delta_{i}+\theta}{\theta+\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}}p_{i}\right|_{\left(\bm{\Lambda}^{(m)},\mathbf{D}\right)},$ (11)
@@ 3.2 EM Algorithm for penMCFM
$El_{c_{1}}^{(m)}(\mathbf{b})=\sum_{i=1}^{n}\left[p_{i}^{(m)}\mathbf{bz}_{i}\mathbf{-}\log\left(1+e^{\mathbf{bz}_{i}}\right)\right]-n\sum_{j=1}^{P^{p}_{1}}P_{\mathbf{\lambda}_{1}}(b_{p,j}),$ (12)
$El^{(m)}_{c_{2}}(\alpha,\gamma,\bm{\beta})=\sum_{i=1}^{n}\left[\delta_{i}\log\left(\alpha\gamma t_{i}^{\gamma-1}e^{\bm{\beta}\mathbf{x}_{i}}\right)-c_{i}^{(m)}\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}\right]-n\sum_{l=1}^{P^{p}_{2}}P_{\mathbf{\lambda}_{2}}(\beta_{p,l}),$ (13)
$El^{(m)}_{c_{3}}(\theta)=\sum_{i=1}^{n}\left[\theta\log\theta-\log\Gamma(\theta)-a_{i}^{(m)}\theta+(\delta_{i}+\theta-1)b_{i}^{(m)}\right].$ (14)
@@
$El_{c_{1}}^{(m)}(\mathbf{b)=}\frac{1}{n}\sum_{i=1}^{n}\left[p_{i}^{(m)}\mathbf{bz}_{i}\mathbf{-}\log\left(1+e^{\mathbf{bz}_{i}}\right)\right]-\lambda_{1,Enet}\left[\frac{(1-\alpha_{Enet})}{2}\sum_{j=1}^{P^{p}_{1}}b_{p,_{j}}^{2}+\alpha_{Enet}\sum_{j=1}^{P^{p}_{1}}w_{j}\left|b_{p,j}\right|\right],$
$El^{(m)}_{c_{2}}(\alpha,\gamma,\bm{\beta})=\frac{1}{n}\sum_{i=1}^{n}\left[\delta{}_{i}\log\left(\alpha\gamma t_{i}^{\gamma-1}e^{\bm{\beta}\mathbf{x}_{i}}\right)-c_{i}^{(m)}\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}\right]-\lambda_{2,Enet}\left[\frac{(1-\alpha_{Enet})}{2}\sum_{l=1}^{P^{p}_{2}}\beta_{p,_{l}}^{2}+\alpha_{Enet}\sum_{l=1}^{P^{p}_{2}}w_{l}\left|\beta_{p,l}\right|\right],$
@@
$w_{j}^{(k)}=\frac{1}{\left|\widehat{b}_{p,j}^{(k-1)}\right|},\text{ }j=1,...,P_{1},\quad w_{l}^{(k)}=\frac{1}{\left|\widehat{\beta}_{p,l}^{(k-1)}\right|},\text{ }l=1,...,P_{2}.$
@@
$-El_{c_{1}}^{(m)}(\mathbf{b)}=-\frac{1}{n}\sum_{i=1}^{n}\left[p_{i}^{(m)}\mathbf{bz}_{i}\mathbf{-}\log\left(1+e^{\mathbf{bz}_{i}}\right)\right]+\lambda_{1,Enet}\left[\frac{(1-\alpha_{Enet})}{2}\sum_{j=1}^{P^{p}_{1}}b_{p,_{j}}^{2}+\alpha_{Enet}\sum_{j=1}^{P^{p}_{1}}w_{j}^{(k)}\left|b_{p,j}\right|\right],$ (15)
$-El^{(m)}_{c_{2}}(\alpha,\gamma,\bm{\beta})=-\frac{1}{n}\sum_{i=1}^{n}\left[\delta_{i}\log\left(\alpha\gamma t_{i}^{\gamma-1}e^{\bm{\beta}\mathbf{x}_{i}}\right)-c_{i}^{(m)}\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}}\right]+\lambda_{2,Enet}\left[\frac{(1-\alpha_{Enet})}{2}\sum_{l=1}^{P^{p}_{2}}\beta_{p,_{l}}^{2}+\alpha_{Enet}\sum_{l=1}^{P^{p}_{2}}w_{l}^{(k)}\left|\beta_{p,l}\right|\right].$ (16)
@@ 3.3 Implementation of the penMCFM Algorithm
At the $m$th iteration of the E-step, $El_{c_{1}}^{(m)}(\mathbf{b)}$ in (12) has the same structure as the penalized log-likelihood function of a logistic regression model, except for the parameters $p_{i}^{(m)}$. As evident from the definition in (8), $p_{i}^{(m)}$ is not constrained to binary values, but can take any value in the interval $\left[0,1\right]$. Notably, it is important to note that these response values $p_{i}^{(m)}$ are not directly observable within this framework, in contrast to the typical situation in penalized logistic regression problems. Consequently, optimizing $El_{c_{1}}^{(m)}(\mathbf{b)}$ poses more challenges compared to solving a conventional penalization problem. This challenge has been recently highlighted by Bussy et al. [15] and Fu et al. [35], and implies the impossibility to directly employ the glmnet R package [34] to address the high-dimensional component of our problem. The second term, $El_{c_{2}}^{(m)}(\alpha,\gamma,\bm{\beta)}$ in (13), can be considered as the log-likelihood function of the penalized Weibull regression model with $h(t_{i})=\alpha\gamma t_{i}^{\gamma-1}e^{\bm{\beta}\mathbf{x}_{i}+\log c_{i}^{(m)}}$ and $S(t_{i})=\exp\left(-\alpha t_{i}^{\gamma}e^{\bm{\beta}\mathbf{x}_{i}+\log c_{i}^{(m)}}\right).$ Here $c_{i}^{(m)}$ is a constant that is evaluated by using (11) before the optimization steps of the algorithm. We employ the lbfgs R package [22], which provides the functionalities to utilize $L_{1}$ penalties for the minimization of the cost functions outlined in (15) and (16). Finally the third term, $El_{c_{3}}^{(m)}(\theta\mathbf{)}$ in (14), is maximized to obtain the estimate of the frailty distribution parameter $\theta$ (or its reciprocal, i.e., the variance).
We give a pseudo-code description of our EM algorithm for penMCFM in Algorithm 1. Initialization is carried out as follows: we use Cox regression estimates for $\bm{\beta}_{u}$, moment estimates for the Weibull distribution parameters $\alpha$ and $\gamma$, and we set $\bm{\beta}_{p}=\mathbf{b}_{0}=\mathbf{b}_{u}=\mathbf{b}_{p}=0$ and $\theta=1$. Generally, it is assumed that the tuning parameters for both penalties are common, i.e. $\bm{\lambda}_{1,Enet}=\bm{\lambda}_{2,Enet}=\bm{\lambda}_{Enet}$ for the sake of simplicity as outlined in [35]. We have followed this approach in our simulations as well. We then generate a sequence of $\bm{\lambda}_{Enet}$ based on the penalized Cox regression of $\mathbf{t},\>\bm{\delta},\>\mathbf{x}_{u},\>\mathbf{x}_{p}$ using the glmnet R package [34].
@@
Initialization: $\bm{\beta}_{u}=\bm{\widehat{\beta}_{Cox}},\>\bm{\beta}_{p}=0,\>\mathbf{b}=0,\>\alpha=\widehat{\alpha}_{ME},\>\gamma=\widehat{\gamma}_{ME},\>\theta=1$ Determine: $\bm{\lambda}_{Enet}$ based on $\mathbf{t},\>\bm{\delta},\>\mathbf{x_{u},\>x_{p}}$ Input: $\bm{\lambda}_{Enet},\>\alpha_{Enet},\>\alpha,\>\gamma,\>\theta,\>\bm{\beta_{u}},\>\bm{\beta_{p}},\>\mathbf{b},\>\epsilon$
@@
             $m\leftarrow++$ E-step: compute $p_{i}^{(m)}$, $a_{i}^{(m)}$, $b_{i}^{(m)}$ and $c_{i}^{(m)}$ from (8)-(11);
             M-step: update $\mathbf{b_{p}}^{(m)}$ minimizing (15) update $\bm{\beta_{p}}^{(m)}$ minimizing (16) update $b_{0}^{(m)}$, $\mathbf{b_{u}}^{(m)}$ minimizing (15) given $\mathbf{b_{p}}^{(m)}$;
             update $\alpha^{(m)}$, $\gamma^{(m)}$ and $\bm{\beta_{u}}^{(m)}$ minimizing (16) given $c_{i}^{(m)}$, $\bm{\beta_{p}}^{(m)}$ update $\bm{\theta}^{(m)}$ minimizing $-El^{(m)}_{c_{3}}(\theta)$ in (14) given $a_{i}^{(m)}$, $b_{i}^{(m)}$
      until  $\left|El_{c_{1}}^{(m)}(\mathbf{b}^{(m)})-El_{c_{1}}^{(m-1)}(\mathbf{b}^{(m-1)})\right|<\epsilon$, $\left|El_{c_{2}}^{(m)}(\alpha^{(m)},\gamma^{(m)},\bm{\beta}^{(m)})-El_{c_{2}}^{(m-1)}(\alpha^{(m-1)},\gamma^{(m-1)},\bm{\beta}^{(m-1)})\right|<\epsilon$, $\left|El_{c_{3}}^{(m)}(\theta^{(m)})-El_{c_{3}}^{(m-1)}(\theta^{(m-1)})\right|<\epsilon$ ;

### v2

@@ 3 Inference in the Penalized Mixture Cure Frailty Model
$L(\alpha,\gamma,\theta,\bm{\beta},\mathbf{b}|\mathbf{D})$ $=$ $\prod_{i=1}^{n}\left\{f_{pop}(t_{i}|\mathbf{x}_{i}\mathbf{,z}_{i})\right\}^{\delta_{i}}\left\{S_{pop}(t_{i}|\mathbf{x}_{i}\mathbf{,z}_{i})\right\}^{1-\delta_{i}}$ (6)
$=$ $\prod_{i=1}^{n}\left\{\frac{e^{\mathbf{z}_{i}^{\top}\mathbf{b}}}{1+e^{\mathbf{z}_{i}^{\top}\mathbf{b}}}\alpha\gamma t_{i}^{\gamma-1}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\left(1+\frac{\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}}{\theta}\right)^{-\theta-1}\right\}^{\delta_{i}}$
×$\times\left\{\frac{1}{1+e^{\mathbf{z}_{i}^{\top}\mathbf{b}}}\left(1+e^{\mathbf{z}_{i}^{\top}\mathbf{b}}\left(1+\frac{\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}}{\theta}\right)^{-\theta}\right)\right\}^{1-\delta_{i}}$
where $\alpha,\gamma$ are the unknown parameters of the Weibull distribution, $\theta$ is the unknown parameter of the gamma distribution used for the frailty, $\mathbf{b}\in\mathbb{R}^{P_{1}+1}$ and $\bm{\beta}\in\mathbb{R}^{P_{2}}$ are vectors of unknown regression coefficients for the covariates $\mathbf{x}$ and $\mathbf{z}$, respectively, where we indicate $\mathbf{z}_{i}^{\top}\mathbf{b}=b_{0}+\mathbf{z}_{u,i}^{\top}\mathbf{b}_{u}+\mathbf{z}_{p,i}^{\top}\mathbf{b}_{p}$ and $\mathbf{x}_{i}^{\top}\bm{\beta}=\mathbf{x}_{u,i}^{\top}\bm{\beta}_{u}+\mathbf{x}_{p,i}^{\top}\bm{\beta}_{p}$, $i=1,...,n$. The maximum likelihood estimator of the unknown parameters can be obtained by direct maximization of the observed likelihood function in (6) when the number of unknown parameters is not large. Since we would like to consider high-dimensional covariates and perform variable selection via penalization, direct maximization of the likelihood function will not provide parameter estimates. We therefore propose to adapt the E-M algorithm to include regularization methods to the purposes of parameter estimation and variable selection in penMCFM. Details on the proposed implementation of the EM for penMCFM are given in Section 3.2, after recalling the standard EM for MCFM in Section 3.1. For the sake of comparisons in the simulation studies, we have also adapted the GMIFS method from [36] to obtain parameter estimations in penMCFM using the observed likelihood function in (6), and details on the implementation of this method are given in Section 3.4.
@@ 3.1 EM Algorithm for the standard MCFM
$S_{pop}(t_{i}|\mathbf{x}_{i}\mathbf{,z}_{i}\mathbf{,}w_{i})=1-\pi(\mathbf{z}_{i})+\pi(\mathbf{z}_{i})\exp\left(-w_{i}\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\right),$
@@
$S(t_{i}|\mathbf{x}_{i}\mathbf{,}w_{i},Y_{i}\mathbf{=}1)=\exp\left(-w_{i}\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\right),$
and
$h(t_{i}|\mathbf{x,}w_{i},Y_{i}\mathbf{=}1)=w_{i}\alpha\gamma t_{i}^{\gamma-1}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}.$
@@
$\pi(\mathbf{z}_{i})S(t_{i}|\mathbf{x}_{i}\mathbf{,}w_{i},Y_{i}\mathbf{=}1)\left[h(t_{i}|\mathbf{x_{i}},w_{i},Y_{i}\mathbf{=}1)\right]^{\delta_{i}}$
@@
$l_{c_{2}}(\alpha,\gamma,\bm{\beta})=\sum_{i=1}^{n}\left[\delta_{i}\log\left(\alpha\gamma t_{i}^{\gamma-1}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\right)-w_{i}y_{i}\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\right],$
@@
$p_{i}^{(m)}\equiv E(y_{i}|\bm{\Lambda}^{(m)},\mathbf{D})=\delta_{i}+(1-\delta_{i})\left.\frac{\pi(\mathbf{z}_{i})\left(1+\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}/\theta\right)^{-\theta}}{(1-\pi(\mathbf{z}_{i}))+\pi(\mathbf{z}_{i})\left(1+\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}/\theta\right)^{-\theta}}\right|_{\left(\bm{\Lambda}^{(m)},\mathbf{D}\right)},$ (8)
$a_{i}^{(m)}\equiv E(w_{i}|\bm{\Lambda}^{(m)},\mathbf{D})=\frac{\delta_{i}+\theta}{\theta+\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}}p_{i}+\left.\frac{\delta_{i}+\theta}{\theta}(1-p_{i})\right|_{\left(\bm{\Lambda}^{(m)},\mathbf{D}\right)},$ (9)
$b_{i}^{(m)}\equiv E\left(\log w_{i}|\bm{\Lambda}^{(m)},\mathbf{D}\right)=\left[\varphi(\delta_{i}+\theta)-\log\theta\right](1-p_{i})+\left.\left[\varphi(\delta_{i}+\theta)-\log(\theta+\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}})\right]p_{i}\right|_{\left(\bm{\Lambda}^{(m)},\mathbf{D}\right)},$ (10)
$c_{i}^{(m)}\equiv E\left(w_{i}y_{i}|\bm{\Lambda}^{(m)},\mathbf{D}\right)=\left.\frac{\delta_{i}+\theta}{\theta+\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}}p_{i}\right|_{\left(\bm{\Lambda}^{(m)},\mathbf{D}\right)},$ (11)
@@ 3.2 EM Algorithm for penMCFM
$El_{c_{1}}^{(m)}(\mathbf{b})=\sum_{i=1}^{n}\left[p_{i}^{(m)}\mathbf{z}_{i}^{\top}\mathbf{b}-\log\left(1+e^{\mathbf{z}_{i}^{\top}\mathbf{b}}\right)\right]-n\sum_{j=1}^{P^{p}_{1}}P_{\mathbf{\lambda}_{1}}(b_{p,j}),$ (12)
$El^{(m)}_{c_{2}}(\alpha,\gamma,\bm{\beta})=\sum_{i=1}^{n}\delta_{i}\log\left(\alpha\gamma t_{i}^{\gamma-1}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\right)-\sum_{i=1}^{n}c_{i}^{(m)}\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}-n\sum_{l=1}^{P^{p}_{2}}P_{\mathbf{\lambda}_{2}}(\beta_{p,l}),$ (13)
$El^{(m)}_{c_{3}}(\theta)=\sum_{i=1}^{n}\left[(\delta_{i}+\theta-1)b_{i}^{(m)}-a_{i}^{(m)}\theta\right]+n(\theta\log\theta-\log\Gamma(\theta)).$ (14)
@@
$El_{c_{1}}^{(m)}(\mathbf{b)}=\frac{1}{n}\sum_{i=1}^{n}\left[p_{i}^{(m)}\mathbf{z}_{i}^{\top}\mathbf{b}-\log\left(1+e^{\mathbf{z}_{i}^{\top}\mathbf{b}}\right)\right]-\lambda_{1,Enet}\left[\frac{(1-\alpha_{Enet})}{2}\sum_{j=1}^{P^{p}_{1}}b_{p,_{j}}^{2}+\alpha_{Enet}\sum_{j=1}^{P^{p}_{1}}w_{j}\left|b_{p,j}\right|\right],$
$El^{(m)}_{c_{2}}(\alpha,\gamma,\bm{\beta})=\frac{1}{n}\sum_{i=1}^{n}\left[\delta_{i}\log\left(\alpha\gamma t_{i}^{\gamma-1}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\right)-c_{i}^{(m)}\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\right]-\lambda_{2,Enet}\left[\frac{(1-\alpha_{Enet})}{2}\sum_{l=1}^{P^{p}_{2}}\beta_{p,_{l}}^{2}+\alpha_{Enet}\sum_{l=1}^{P^{p}_{2}}w_{l}\left|\beta_{p,l}\right|\right],$
@@
$w_{j}^{(k)}=\frac{1}{\left|\widehat{b}_{p,j}^{(k-1)}\right|},\text{}j=1,...,P_{1},$
and
$w_{l}^{(k)}=\frac{1}{\left|\widehat{\beta}_{p,l}^{(k-1)}\right|},\text{}l=1,...,P_{2},$
@@
$-El_{c_{1}}^{(m)}(\mathbf{b)}=-\frac{1}{n}\sum_{i=1}^{n}\left[p_{i}^{(m)}\mathbf{z}_{i}^{\top}\mathbf{b}-\log\left(1+e^{\mathbf{z}_{i}^{\top}\mathbf{b}}\right)\right]+\lambda_{1,Enet}\left[\frac{(1-\alpha_{Enet})}{2}\sum_{j=1}^{P^{p}_{1}}b_{p,_{j}}^{2}+\alpha_{Enet}\sum_{j=1}^{P^{p}_{1}}w_{j}^{(k)}\left|b_{p,j}\right|\right],$ (15)
$-El^{(m)}_{c_{2}}(\alpha,\gamma,\bm{\beta})$ $=-\frac{1}{n}\sum_{i=1}^{n}\left[\delta_{i}\log\left(\alpha\gamma t_{i}^{\gamma-1}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\right)-c_{i}^{(m)}\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}}\right]$
$+\lambda_{2,Enet}\left[\frac{(1-\alpha_{Enet})}{2}\sum_{l=1}^{P^{p}_{2}}\beta_{p,_{l}}^{2}+\alpha_{Enet}\sum_{l=1}^{P^{p}_{2}}w_{l}^{(k)}\left|\beta_{p,l}\right|\right].$ (16)
@@ 3.3 Implementation of the penMCFM Algorithm
At the $m$th iteration of the E-step, $El_{c_{1}}^{(m)}(\mathbf{b)}$ in (12) has the same structure as the penalized log-likelihood function of a logistic regression model, except for the parameters $p_{i}^{(m)}$. As evident from the definition in (8), $p_{i}^{(m)}$ is not constrained to binary values, but can take any value in the interval $\left[0,1\right]$. Notably, it is important to note that these response values $p_{i}^{(m)}$ are not directly observable within this framework, in contrast to the typical situation in penalized logistic regression problems. Consequently, optimizing $El_{c_{1}}^{(m)}(\mathbf{b)}$ poses more challenges compared to solving a conventional penalization problem. This challenge has been recently highlighted by Bussy et al. [31] and Fu et al. [35], and implies the impossibility to directly employ the glmnet R package [16] to address the high-dimensional component of our problem. The second term, $El_{c_{2}}^{(m)}(\alpha,\gamma,\bm{\beta)}$ in (13), can be considered as the log-likelihood function of the penalized Weibull regression model with $h(t_{i})=\alpha\gamma t_{i}^{\gamma-1}e^{\mathbf{x}_{i}^{\top}\bm{\beta}+\log c_{i}^{(m)}}$ and $S(t_{i})=\exp\left(-\alpha t_{i}^{\gamma}e^{\mathbf{x}_{i}^{\top}\bm{\beta}+\log c_{i}^{(m)}}\right).$ Here $c_{i}^{(m)}$ is a constant that is evaluated by using (11) before the optimization steps of the algorithm. We employ the lbfgs R package [38], which provides the functionalities to utilize $L_{1}$ penalties for the minimization of the cost functions outlined in (15) and (16). Finally the third term, $El_{c_{3}}^{(m)}(\theta\mathbf{)}$ in (14), is maximized to obtain the estimate of the frailty distribution parameter $\theta$ (or its reciprocal, i.e., the variance).
We give a pseudo-code description of our EM algorithm for penMCFM in Algorithm 1. Initialization is carried out as follows: we use Cox regression estimates for $\bm{\beta}_{u}$, moment estimates for the Weibull distribution parameters $\alpha$ and $\gamma$, and we set $\bm{\beta}_{p}=\mathbf{b}_{0}=\mathbf{b}_{u}=\mathbf{b}_{p}=0$ and $\theta=1$. Generally, it is assumed that the tuning parameters for both penalties are common, i.e. $\bm{\lambda}_{1,Enet}=\bm{\lambda}_{2,Enet}=\bm{\lambda}_{Enet}$ for the sake of simplicity. We have followed this approach in our simulations as well. We then generate a sequence of $\bm{\lambda}_{Enet}$ based on the penalized Cox regression of $\mathbf{t},\>\bm{\delta},\>\mathbf{x}_{u},\>\mathbf{x}_{p}$ using the glmnet R package [16].
@@
[…]
γ=^γME,θ=1
Determine: $\bm{\lambda}_{Enet}$ based on $\mathbf{t},\>\bm{\delta},\>\mathbf{x_{u},\>x_{p}}$
Input: $\bm{\lambda}_{Enet},\>\alpha_{Enet},\>\alpha,\>\gamma,\>\theta,\>\bm{\beta_{u}},\>\bm{\beta_{p}},\>\mathbf{b},\>\epsilon$
@@
            $m\leftarrow++$
            E-step:
            compute $p_{i}^{(m)}$, $a_{i}^{(m)}$, $b_{i}^{(m)}$ and $c_{i}^{(m)}$ from (8)-(11);
            M-step:
            update $\mathbf{b_{p}}^{(m)}$ minimizing (15)
            update $\bm{\beta_{p}}^{(m)}$ minimizing (16)
            update $b_{0}^{(m)}$, $\mathbf{b_{u}}^{(m)}$ minimizing (15) given $\mathbf{b_{p}}^{(m)}$;
            update $\alpha^{(m)}$, $\gamma^{(m)}$ and $\bm{\beta_{u}}^{(m)}$ minimizing (16) given $c_{i}^{(m)}$, $\bm{\beta_{p}}^{(m)}$
            update $\bm{\theta}^{(m)}$ minimizing $-El^{(m)}_{c_{3}}(\theta)$ in (14) given $a_{i}^{(m)}$, $b_{i}^{(m)}$
      until  $\left|El_{c_{1}}^{(m)}(\mathbf{b}^{(m)})-El_{c_{1}}^{(m-1)}(\mathbf{b}^{(m-1)})\right|<\epsilon$ ,
$\left|El_{c_{2}}^{(m)}(\alpha^{(m)},\gamma^{(m)},\bm{\beta}^{(m)})-El_{c_{2}}^{(m-1)}(\alpha^{(m-1)},\gamma^{(m-1)},\bm{\beta}^{(m-1)})\right|<\epsilon$
$\left|El_{c_{3}}^{(m)}(\theta^{(m)})-El_{c_{3}}^{(m-1)}(\theta^{(m-1)})\right|<\epsilon$